OpenStudy (anonymous):

Simplify the sum. (w^2 + 2w - 24/ w^2 + w - 30 ) + 8/ w - 5

OpenStudy (anonymous):

First factor the top and bottom of the fraction

OpenStudy (anonymous):

(w + 12)(w - 2)/(w - 5)(w + 6) @BluFoot

OpenStudy (anonymous):

\[\frac{ (w + 12)(w - 2) }{ ( w - 5)(w + 6)} + \frac{ 8 }{ (w - 5) }\]

OpenStudy (amoodarya):

you have a mistake (w+12)(w-2)=w2+10w-24 look the image

OpenStudy (anonymous):

( w + 6) ( w - 4) @amoodarya

OpenStudy (amoodarya):

i write the wrong expression ! did you look th efile?